# PIMRL: Physics-Informed Multi-Scale Recurrent Learning for Spatiotemporal Prediction

### v2

@@ 4 Experiments
Datasets. We conducted experiments on five datasets, including: Korteweg-de Vries (KdV), 2D Burgers, FitzHugh-Nagumo (FN), 2D Gray-Scott (2D GS), and 3D Gray-Scott (3D GS). The IC for the Kdv equation is created by summing multiple sine waves with random amplitudes, phases, and frequencies, resulting in a complex waveform. ICs for the Burgers’ equation are generated randomly according to a Gaussian distribution. The FN equation is initialized with random Gaussian noise for a warm-up period, after which time sequences are extracted to form the dataset. The GS equation starts the reaction from random initial positions and then diffuses. In these cases, except for the Kdv case, which was solved using the Finite Volume Method (FVM), the rest were solved using the Finite Difference Method (FDM). Additionally, We have two sets of data with different time scales originating from the same ICs. The micro-scale data $\textbf{U}^{\text{micro}}=\{\textbf{u}_{0},\textbf{u}_{\delta t},\textbf{u}_{2\delta t}\cdot\cdot\cdot\cdot\}\in\text{R}^{\text{micro}}$ is characterized by short and scattered continuous time intervals, while the macro-scale data $\textbf{U}^{\text{macro}}=\{\textbf{u}_{0},\textbf{u}_{\Delta t},\textbf{u}_{2\Delta t}\cdot\cdot\cdot\cdot\textbf{u}_{\text{Tend}}\}\in\text{R}^{\text{macro}}$ exhibits persistent continuity until the end. The validation and test sets are established based on different ICs but with the same parameters, making it more challenging than extrapolation under the same ICs. The dataset configuration is presented in Appendix.
Figure 5: An overview of the comparison between our PIMRL framework and baseline models is provided, including error propagation curves (left), error boxplots (middle), and final prediction plots (right). Figures (a) through (e) respectively showcase the qualitative results for the KdV, Burgers, 2D GS, FN, and 3D GS cases.
Model training. The primary objective is to first pretrain the micro-scale module using the micro-scale data to learn the underlying physics. Subsequently, the pretrained micro-scale module is integrated into the overall framework, and the entire model is trained using macro-scale data to capture the spatial evolution patterns over long time stepping. Both the pretraining and training processes can be formulated as auto-regressive rollout problems. The loss function is defined as $L(\boldsymbol{\theta})=\frac{1}{BN}\sum_{i=1}^{B}\sum_{j=1}^{N}(y_{i,j}-\hat{y}_{i,j})^{2}$, where the $B$ and $N$ denote the number of batches and the batch size. The $\boldsymbol{\theta}$ indicates all the trainable parameters and $(y_{i,j}-\hat{y}_{i,j})$ means the difference between the rollout-prediction $\hat{y}_{i,j}$ of $j$-th sample in the $i$-th batch and the corresponding labeled data $y_{i,j}$. More training details are shown in Appendix.
@@
Baseline models. To validate the effectiveness of the proposed PIMRL framework, we introduced several baseline models. Firstly, we considered the widely recognized high-performing data-driven model FNO [10], which has been trained on datasets with two different time intervals, denoted as FNO (trained on fine-scale data with small time steps) and FNO-coarse (trained on coarse-scale data with large time steps). In particular, we examined the impact of cumulative error on long-term predictions. Our results show that the influence of error accumulation is significant. Detailed analysis is provided in Section 4.1. Secondly, we included the PeRCNN model [19], which embeds physical knowledge in a hard way and demonstrates excellent performance on multiple datasets. Due to constraints on the time stepping of the model, PeRCNN is trained only on datasets with small time stepping. Lastly, we incorporated ConvLSTM [35], a classic sequential prediction model trained on datasets with large time stepping, serving as the macro-scale module within our framework in a residual way. The details are shown in the Appendix.
@@
Evaluation Metrics. To comprehensively evaluate the performance of our model, we adopted several metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and High Correction Time (HCT). These metrics provide a multi-faceted assessment of the model’s accuracy and reliability. RMSE is calculated on the macro-scale data to facilitate comparisons between models operating at different granularities. MAE provides a measure of the average absolute difference between the predicted and actual values,less sensitive to outliers compared to RMSE. HCT evaluates the time it takes for the model to correct its predictions to a high level of accuracy, which is particularly important for long-term prediction. Detailed formulas for these metrics are provided in Appendix.
4.1 Main Results
Figure 5 illustrates the performance results of our framework in comparison with various baseline models across different cases, including error propagation, showcasing box plots of different quantities and their prediction outcomes. Additionally, Table 1 presents relevant quantitative metrics as an expression of the results, where * in Table 1 indicates that the inference process has reached the end of the test data.
Table 1: Quantitative results of our model and baselines.
@@ 4.1 Main Results
Promotion ($\uparrow$) 51.5% 35.5% 54.0%
Burgers FNO 0.1561 0.1301 0.104
@@
Promotion ($\uparrow$) 28.4% 80.6% 43.4%
KdV Equation. The KdV equation describes the evolution of nonlinear wave phenomena. In this study, due to the absence of one-dimensional instances in the PeRCNN model, we designed our own physical embedding module and corresponding model, following the principles of PeRCNN as detailed in Appendix. We observed significant discrepancies between the predictions of the PeRCNN model and the ground truth, as shown in the blue box of Figure 5(a). Meanwhile, the other baseline models performed poorly at the start, likely due to the complexity of the KdV equation. In contrast, our proposed model demonstrated substantial advantages over the baseline models in terms of predictive accuracy. It maintained a consistent basic shape with the ground truth values, even in long-term forecasts. As shown in Table 1, the PIMRL framework achieved 30% to 50% improvements in the evaluation metrics, highlighting a significant advancement in the field.
2D Burgers Equation. On the right side of Figure 5(b), we observe that besides our framework, only PeRCNN achieved satisfactory results after training on the micro-scale, while the remaining baseline models failed to capture the physical changes effectively. Additionally, in the areas marked by the blue box in Figure 5(b), we can see errors in PeRCNN’s predictions for long-term forecasts, whereas our PIMRL framework demonstrates superior performance, as also visually depicted in Figure 5(b). In the error propagation shown in Figure 5(b), we notice that although initially our framework exhibits higher RMSE compared to the PeRCNN model due to the macro-scale module, the accumulated errors lead to inaccurate predictions by PeRCNN over time. While PeRCNN and PIMRL exhibit the same performance on the HCT metric, this similarity arises from insignificant changes in the later stages. The cumulative error of PeRCNN does not reflect in this particular metric. As shown in Table 1, there are significant improvements in both RMSE and MAE. Additionally, the noticeable divergence of PeRCNN can be observed in the snapshot of the final step.
2D FitzHugh-Nagumo Equation. In the Figure 5(c), we can observe that the purely data-driven method mentioned earlier fails to fully exhibit its original performance on this relatively small dataset. FNO and FNO-coarse represent the effectiveness of the FNO method when trained on datasets of different granularities, respectively. It is worth noting that the performance of FNO-coarse surpasses that of the FNO model trained on micro-scale data. This result effectively validates our hypothesis that error accumulation can significantly impact the model’s performance in long-term multi-step predictions. Quantitatively, our model has improved by at least 10% compared to the currently best-performing model. Furthermore, in the Table 1 we observe that the variation curves of RMSE and other metrics for our model and the PeRCNN model exhibit a striking similarity, yet our model outperforms PeRCNN. This not only indicates that PeRCNN effectively leverages its supervised refinement capabilities but also demonstrates its ability to mitigate the influence of cumulative errors through the evolution of larger time stepping.
@@
2D Gray-Scott Equation. As shown in the left part of Figure 5(d), only PIMRL effectively carried out long-term predictions, showcasing a clear demonstration of the cumulative errors of PeRCNN in this case. The model exhibited significant deviations from ground truth values over extended periods, highlighting its limitations in capturing long-term variations. Among the other models, FNO-Coarse exhibited the best performance. It is evident that FNO performed poorly in the absence of physically informed embeddings, particularly with small time stepping. In this case, Table 1 further validates the superior performance of PIMRL with both RMSE and MAE showing an improvement of over 70%, and the HCT similarity of PIMRL still at 0.99 at the final prediction step. This underscores the effectiveness of the PIMRL framework in enhancing predictive accuracy and maintaining consistency in the predictions throughout the forecasting process. Such robust performance metrics highlight the potential of PIMRL in addressing complex data-driven challenges.
@@
3D Gray-Scott Equation. As shown on the right side of Figure 5(e), the predictions under our framework closely align with the ground truth. Through box plots and error evolution graphs, it is evident that PeRCNN, as the best-performing model among the baselines in our study, outperformed the other models lacking physical knowledge embeddings, especially when compared to FNO trained on the same dataset. In quantitative analysis of Table 1, the improvements are also significant, not only in the overall evaluation metrics of RMSE and MAE in 28.4% and 80.6% but also in the substantial growth of the HCT in 43.4%.
Table 2: Results for ablation study.
@@
PeRCNN w/o Physics-based FD Conv NaN
4.2 Ablation Study
@@
To evaluate the impact of components of PIMRL and demonstrate the effectiveness of our framework structure, we have designed five novel models and provided their RMSE results in the FN case. (1) The ablation study with the “PIMRL w/o Connect”, where the connections between the micro-module and the macro-module are removed, is designed to demonstrate the effectiveness of the PIMRL framework’s structural design. This experiment, which leaves only the serial structure of the two modules, shows that the connections within the PIMRL framework are essential for its performance. (2) The “FNO-MRL” replaces the micro-scale module containing physical information, PeRCNN, with the data-driven model FNO, aiming to validate the efficacy of the physical embedding in the micro-scale module within our framework. (3) “PIMRL w/o Pretraining” provides a perspective on the training method by eliminating the pretraining step for the micro module. In this ablation study, the absence of pretraining leads to inferior performance compared to the PIMRL. This demonstrates that directly introducing large time intervals for training can deprive the micro module of the opportunity to learn fine-grained changes, similar to why PeRCNN cannot be directly used with large time intervals. (4) “PIMRL w/o Physics-based FD Conv” indicates the removal of the Physics-based FD Convolution. This ablation study emphasizes the effectiveness of the Physics-based FD Conv by showing the performance degradation when it is omitted. (5) “PeRCNN w/o Physics-based FD Conv” is a version of PeRCNN without the Physics-based Finite Difference (FD) Convolution. While this version can initially make relatively accurate predictions, the errors accumulate over subsequent iterations, eventually becoming unacceptably large.
The five ablation studies conducted not only validate the effectiveness of the interaction design by removing the connection mechanisms but also highlight the contribution of our proposed connection approach. Additionally, by substituting the micro-modules with non-physics-embedded data-driven models, the experiments confirm the efficacy of the physics-embedding within our PIMRL framework. Subsequent pre-training ablation experiments and the removal of the Physics-based FD Conv further substantiate the effectiveness of the corresponding methods and modules.
4.3 Inference Time
Figure 6: Computational time for comparison
In the aforementioned cases, PIMRL not only achieves state-of-the-art performance in long-term predictions but also significantly reduces the computational time cost. Compared to traditional methods such as Direct Numerical Simulation (DNS), our framework is substantially faster, demonstrating a significant improvement in computational efficiency (given the same computing facility, aka, CPU). Additionally, PIMRL not only delivers superior predictive accuracy, but also outperforms PeRCNN in terms of computational efficiency. The numerical methods used for DNS are consistent with the parameter settings during data generation. The difference lies in the required prediction time length, which needs to be the same as that of PIMRL and PeRCNN for a fair comparison.
5 Conclusions
This paper introduces a new multi-scale temporal model named Physics-Informed Multi-Scale Recurrent Learning (PIMRL) for prediction of spatiotemporal dynamics. Adhering to the idea of controlling cumulative error by reducing iterations, PIMRL integrates modules across different scales within its framework to realize this concept, and it can efficiently leverage multi-scale data, which facilitates learning physical laws with limited resources. From fluid dynamics to reaction-diffusion systems, PIMRL demonstrates superior performance in handling multi-scale data, providing accurate long rollout predictions. Our future work will optimize the macro-scale model for time-series tasks to improve computational efficiency and predictive accuracy. We also plan to integrate super-resolution techniques to enhance the model’s adaptability to various spatiotemporal scales, broadening its application to more complex physical systems.
Acknowledgement
The work is supported by the National Natural Science Foundation of China (No. 62276269 and No. 92270118), the Beijing Natural Science Foundation (No. 1232009), and the Fundamental Research Funds for the Central Universities (No. 202230265).
APPENDIX
Appendix A Impact statement
@@ Appendix B Physical Filter
$f_{\text{xxx}}=\frac{-f_{i-2}+2f_{i-1}+0f_{i}-2f_{i+1}+f_{i+2}}{2h^{3}}$ (S1)
Figure S1: Filter for KdV
As shown in the Figure S1, we designed a Physical-Filter to represent $\frac{\partial^{3}u}{\partial x^{3}}$ in the KdV equation. This approach leverages the inherent physical properties of the system to accurately model the third-order spatial derivative, thereby enhancing the accuracy and efficiency of the numerical solution. Among the parameters, $h$ indicates the $\Delta x$ in the cases. The Boundary Padding is an approach to adapt to periodic boundary conditions by replacing the original zero padding with a periodic boundary padding.
Appendix C Implementation Details
C.1 Overview
@@
Figure S2: Overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
@@ C.2 Encoder and Decoder
C.3 ResidualLSTM
The structure of ResidualLSTM has been clearly illustrated in the main text. Here, we will elucidate the architecture of the ConvLSTMcell as the Figure S3.
Figure S3: ConvLSTMCell
$c_{i}$ $=\sigma(\text{W}_{xi}(x_{t})+\text{W}_{hi}(h_{t-1}))$ (S2)
@@ C.3 ResidualLSTM
The set of equations S2 presented above outlines the operations of a ConvLSTM cell. The equations involve computations of various gates and states within the cell, including input gate $c_{i}$, forget gate $c_{f}$, cell state $c_{c}$, output $C_{t}$, and the updated hidden state $h_{t}$. These equations govern the flow of information and transformations within the ConvLSTM cell, enabling the model to process spatiotemporal data efficiently by considering spatial dimensions in the calculations. The parameters $\text{W}{xi}$, $\text{W}{hi}$, $\text{W}{xf}$, $\text{W}{hf}$, $\text{W}{xo}$, and $\text{W}{ho}$, depicted in Figure S3, manage information from inputs ($x_{t}$) and history ($h_{t-1}$,$C_{t-1}$) in a convolutional manner.
Appendix D Baseline Models
In order to compare and evaluate the performance of our proposed method, we have trained multiple state-of-the-art (SOTA) baseline models as well as classical models, and compared them with our model. The introductions to each baseline model are presented below, while the training details are outlined in subsequent sections E.
Fourier Neural Operator (FNO) [10]. FNO is a method that combines Fourier transforms with neural networks. This approach comprises two main components. The first component involves performing Fourier transforms on the system state quantities, learning certain information in the frequency domain, and then applying an inverse transform. The second component utilizes convolutions to process the system state quantities, complementing the information not captured during the frequency domain learning. The combination of these two components serves as the final result. We make two FNO models sharing the same architecture, which training in micro and macro scale datasets respectively and inferring in micro and macro time intervals. In the same model, we conducted training separately for two types of data, resulting in FNO and FNO-coarse.
ConvLSTM [35]. ConvLSTM is a specialized neural network architecture that combines convolutional and LSTM layers to effectively model spatial and temporal dependencies in sequential data.
PeRCNN [19]. PeRCNN represents a physics-informed learning methodology, embedding physical laws directly into the neural network architecture. It incorporates multiple parallel convolutional neural networks (CNNs), leveraging the simulation of polynomial equations through feature map multiplication. By doing so, PeRCNN augments the model’s extrapolation and generalization capabilities.
Appendix E Training Details
@@
Appendix F Evaluation Metrics
In this paper, we have adopted some classical evaluation metrics such as RMSE, MAE and HCT. Root Mean Square Error (RMSE) quantifies the average error magnitude between estimated and actual values, serving as a gauge of the model’s precision. Conversely, Mean Absolute Error (MAE) assesses the average absolute disparity between anticipated and observed values, denoting the true scale of discrepancies.
@@ Appendix F Evaluation Metrics
RMSE (Root Mean Square Error): $\quad\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (S3)
@@
HCT (High Correction Time): $\sum_{i=1}^{N}\Delta t\cdot 1(\text{PCC}(y_{i},\tilde{y}_{i})>0.8)$
In the above equationsS3, $n$ represents the number of trajectories, $y_{i}$ represents the true value, and $\hat{y}_{i}$ represents the predicted value of the model. The PCC is the Pearson correlation coefficient, which is a statistical metric used to measure the linear correlation between two variables.
Appendix G Dataset Informations
@@
Table S1: Summary of experimental settings for different cases.(The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training)
@@
$\frac{\partial u}{\partial t}=-u\frac{\partial u}{\partial x}-\frac{\partial^{3}u}{\partial x^{3}}$ (S4)
@@
$\frac{\partial u}{\partial t}$ $=-uu_{x}-vu_{y}+\nu\nabla^{2}u,$ (S5)
$\frac{\partial v}{\partial t}$ $=-uv_{x}-vv_{y}+\nu\nabla^{2}v.$ (S6)
@@
$\frac{\partial u}{\partial t}$ $=\mu_{u}\nabla^{2}u+u-u^{3}-v+\alpha,$ (S7)
$\frac{\partial v}{\partial t}$ $=\mu_{v}\nabla^{2}v+(u-v)\beta.$ (S8)
@@
$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+F(1-u),$ (S9)
$\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(F+k)v.$ (S10)
@@
In the three-dimensional case, we have the parameters: $DA=0.2$, $DB=0.1$, $F=0.025$, and $k=0.055$. We got 3 trajectories for training, 1 trajectory for validation and 2 trajectories for testing, where $\Delta t=15\delta t$ and $\delta t=0.25s$.
Appendix H Supplement results
Considering the constraints on the length of the main text, we have placed some important supplementary results in the appendix of the article. First, we conducted additional experiments on the FN equation, expanding our tests to include more advanced baselines such as F-FNO, U-NO, and MWT (shown in table S2). The experimental results indicate that data-driven methods do not perform well on small training datasets, which further corroborates the effectiveness of the PIMRL proposed in our paper.
Table S2: The Result of U-NO, F-FNO, MWT and FNO in the FN Case.
Metrics U-NO F-FNO MWT FNO
RMSE 0.3675 0.2280 0.3494 0.1878
MAE 0.1465 0.1350 0.2228 0.1634
Table S3: Results with Error Bar under RMSE metric on all Cases.
Model PIMRL PeRCNN
Kdv $0.0457\pm 0.0053$ $0.0942\pm 0.0082$
Burgers $0.0068\pm 0.0006$ $0.0075\pm 0.0008$
2DGS $0.0133\pm 2.4\times 10^{-12}$ $0.0455\pm 1.9\times 10^{-11}$
FN $0.1349\pm 0.0040$ $0.1591\pm 0.0061$
3DGS $0.0381\pm 0.0015$ $0.0532\pm 0.0027$
Table S4: The experiments between PIMRL and baseline models on the Burgers example of FNO benchmark.

### v4

@@ Introduction
Datasets. We conducted experiments on five datasets, including Korteweg-de Vries (KdV), Burgers, FitzHugh-Nagumo (FN), 2D Gray-Scott (2D GS), and 3D Gray-Scott (3D GS). The IC for the KdV equation is created by summing multiple sine waves with random amplitudes, phases, and frequencies, resulting in a complex waveform. ICs for the Burgers’ equation are generated randomly according to a Gaussian distribution. The FN equation is initialized with random Gaussian noise for a warm-up period, after which time sequences are extracted to form the dataset. The GS equation starts the reaction from random initial positions and then diffuses. In these cases, except for the KdV case, which was solved using the Finite Volume Method (FVM), the rest were solved using the Finite Difference Method (FDM).
Additionally, we have two sets of data with different time scales originating from the same ICs. The micro-scale data $\mathbf{U}^{\text{micro}}=\{\mathbf{u}_{0},\mathbf{u}_{\delta t},\mathbf{u}_{2\delta t}\cdot\cdot\cdot\}\in\mathbf{R}^{\text{micro}}$ is characterized by short and scattered continuous time intervals, while the macro-scale data $\mathbf{U}^{\text{macro}}=\{\mathbf{u}_{0},\mathbf{u}_{\Delta t},\mathbf{u}_{2\Delta t}\cdot\cdot\cdot\mathbf{u}_{T_{\text{end}}}\}\in\mathbf{R}^{\text{macro}}$ exhibits persistent continuity until the end. The validation and test sets are established based on different ICs but with the same parameters, making it more challenging than extrapolation under the same ICs.
@@
Baseline Models. To validate the effectiveness of the proposed PIMRL framework, we introduced several baseline models. Firstly, we considered the widely recognized high-performing data-driven model FNO (li2020fourier), which has been trained on datasets with two different time intervals, denoted as FNO (trained on fine-scale data with small time steps) and FNO-coarse (trained on coarse-scale data with large time steps). Secondly, we included the PeRCNN model (Rao_2023), which embeds physical knowledge in a hard way. Due to constraints on the time stepping of the model, PeRCNN is trained on datasets with small time steps. Lastly, we incorporated CROP (gao2025discretizationinvariance)(CROP-coarse trained on coarse-scale data with large time steps), a method proposed to address problems with generalization and discretization mismatch errors in existing neural operators across different data resolutions. Since the original CROP paper did not include 1D and 3D cases, we employ AE-ConvLSTM (vlachas2022multiscale) as a substitute. The details are shown in Appendix C.
@@
Evaluation Metrics. To comprehensively evaluate the performance of our model, we adopted metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and High Correction Time (HCT). RMSE is calculated on the macro-scale data to facilitate comparisons between models operating at different granularities. MAE provides a measure of the average absolute difference between the predicted and actual values. HCT evaluates the time it takes for the model to correct its predictions to a high level of accuracy. Detailed formulas for these metrics are provided in the Evaluation Metrics part of Appendix E.
@@
Promotion 51.5% 35.5% 54.0%
Burgers CROP-coarse 0.1103 0.0887 0.064
FNO 0.1561 0.1301 0.104
@@
Promotion 28.4% 80.6% 43.4%
Table 2: Quantitative results of our model and baselines, where * denotes that the prediction lasts to the end.
Main Results
Figure 5 compares our framework with baseline models across cases using box plots of predicted quantities. Table 2 reports the corresponding quantitative metrics.
1D KdV Equation. The KdV equation governs nonlinear wave evolution. Notably, significant discrepancies between PeRCNN predictions and ground truth emerge in the blue box of Figure 4, whereas other baselines exhibit poor initial performance attributable to KdV complexity. Conversely, PIMRL demonstrates substantial advantages in predictive accuracy over baseline models, achieving 30%–50% improvements in evaluation metrics (Table 2) and thereby highlighting significant field advancement.
2D Burgers Equation. Figure 5(a) demonstrates that only PeRCNN and our framework attain satisfactory post-training results on micro-scale data, while other baselines fail to capture physical changes. In the blue-boxed regions, PeRCNN shows significant long-term forecast errors compared to PIMRL’s sustained accuracy. Corresponding quantitative improvements in RMSE and MAE are documented in Table 2, with PeRCNN’s divergence further confirmed at the final time step.
2D FitzHugh-Nagumo Equation. In Figure 5(b), the purely data-driven method fails to achieve its original performance on this relatively small dataset. Quantitatively, our model outperforms the best existing model by at least 10%. Additionally, Table 2 shows that while the RMSE and other metric curves of our model and PeRCNN are similar, our model consistently achieves better results.
@@
2D Gray-Scott Equation. As shown in the left part of Figure 5(c), only PIMRL effectively carried out long-term predictions, showcasing a clear demonstration of the cumulative errors of PeRCNN in this case. In this case, Table 2 further validates the superior performance of PIMRL with both RMSE and MAE showing an improvement of over 70%, and the HCT similarity of PIMRL at 0.99 at the final prediction step. This highlights the effectiveness of PIMRL in improving predictive accuracy and consistency throughout forecasting, demonstrating its potential for tackling complex challenges.
Figure 6: Error propagation curves and final prediction plots for the 3D GS, comparing the PIMRL and baseline models.
@@
3D Gray-Scott Equation. As shown on the right side of Figure 6, the predictions under our framework closely align with the ground truth. Through box plots and error evolution graphs, it is evident that PeRCNN, as the best-performing model among the baselines in our study, outperformed the other models lacking physical knowledge embeddings, especially when compared to FNO trained on the same dataset. In quantitative analysis of Table 2, the improvements are also significant, not only in the overall evaluation metrics of RMSE and MAE in 28.4% and 80.6% but also in the substantial growth of the HCT in 43.4%.
Ablation Study
We designed five novel models and report them in Table 3.
@@
PIMRL 0.1349
Table 3: Results for ablation study in FN-case.
@@
(1) The ablation study with the “PIMRL w/o Connect”, where the connections between the micro-module and the macro-module are removed, is designed to demonstrate the effectiveness of the PIMRL framework’s structural design. This experiment, which leaves only the serial structure of the two modules, shows that the connections within the PIMRL framework are essential. (2) The “FNO-MRL” replaces the micro-scale module containing physical information, PeRCNN, with the data-driven model FNO, aiming to validate the efficacy of the physical embedding in the micro-scale module within our framework. (3) “PIMRL w/o Pretraining” provides a perspective on the training method by eliminating the pretraining step for the micro module. In this ablation study, the absence of pretraining leads to inferior performance compared to the PIMRL. This demonstrates that directly introducing large time intervals for training can deprive the micro module of the opportunity to learn fine-grained changes, similar to why PeRCNN cannot be directly used with large time intervals. (4) “PIMRL w/o Physics-based FD Conv” indicates the removal of the Physics-based FD Convolution. This ablation study emphasizes the effectiveness of the Physics-based FD Conv by showing the performance degradation when it is omitted. (5) “PeRCNN w/o Physics-based FD Conv” is a version of PeRCNN without the Physics-based Finite Difference (FD) Convolution.
Addition Results
To comprehensively evaluate PIMRL, we conducted a series of supplementary experiments, the details of which are provided in Appendix G. In particular, we assessed computational efficiency by comparing runtime, parameter count, and memory usage with other physics-embedded methods, and examined the model’s scalability and practicality in various scenarios. Furthermore, we investigated the robustness of PIMRL under different training set sizes and analyzed the reliability of its predictions using error bar analyses. These additional experiments collectively demonstrate the promising efficiency, scalability, and robustness of PIMRL.
Conclusion
In this paper, we introduced the Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework, the first, to our knowledge, specifically designed to predict the evolution of spatiotemporal systems from multi-scale, burst-sampled data. Its novel architecture synergistically combines a macro-scale module for efficient long-term forecasting with a physics-informed micro-scale module for periodic correction, a design that effectively mitigates cumulative error while capturing underlying physical laws from extremely limited observations. Rigorous evaluations confirmed its substantial performance gains over state-of-the-art baselines, including a reduction in RMSE of up to 70% on the 2D Gray-Scott dataset, thereby validating the effectiveness of this multi-scale, physics-informed approach.
While we acknowledge the framework’s current limitations, such as the yet unexplored sensitivity to the duration of data gaps and the presence of input noise, as detailed in Appendix H, we view them as clear and important avenues for future research. Going forward, we will improve the macro-module’s efficiency and integrate advanced super-resolution to better exploit the model’s potential, enhance generalization, and widen its impact in scientific computing.
Acknowledgments
The work is supported by the Beijing Natural Science Foundation (No. 1232009) and the National Natural Science Foundation of China (No. 62276269 and No. 62506367). R.Z. would like to acknowledge the supported by the China Postdoctoral Science Foundation under Grant Number 2025M771582 and the Postdoctoral Fellowship Program of CPSF under Grant Number GZB20250408. Additional experimental details and full appendices are available in the arXiv version of this paper at https://arxiv.org/abs/2503.10253.
Impact statement
@@ Acknowledgments
$f_{\text{xxx}}=\frac{-f_{i-2}+2f_{i-1}+0f_{i}-2f_{i+1}+f_{i+2}}{2h^{3}}$ (6)
As shown in the Figure LABEL:filter4kdv, we designed a Physical-Filter to represent $\frac{\partial^{3}u}{\partial x^{3}}$ in the KdV equation. This approach leverages the inherent physical properties of the system to accurately model the third-order spatial derivative, thereby enhancing the accuracy and efficiency of the numerical solution. Among the parameters, $h$ indicates the $\Delta x$ in the cases. The Boundary Padding is an approach to adapt to periodic boundary conditions by replacing the original zero padding with a periodic boundary padding.
B: Implementation Details
Overview
@@
Figure 7: Overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
@@
ResidualLSTM
The structure of ResidualLSTM has been clearly illustrated in the main text. Here, we will elucidate the architecture of the ConvLSTMcell as the Figure 8.
Figure 8: ConvLSTMCell
$c_{i}$ $=\sigma(\text{W}_{xi}(x_{t})+\text{W}_{hi}(h_{t-1}))$ (7)
@@
The set of equations 7 presented above outlines the operations of a ConvLSTM cell. The equations involve computations of various gates and states within the cell, including input gate $c_{i}$, forget gate $c_{f}$, cell state $c_{c}$, output $C_{t}$, and the updated hidden state $h_{t}$. These equations govern the flow of information and transformations within the ConvLSTM cell, enabling the model to process spatiotemporal data efficiently by considering spatial dimensions in the calculations. The parameters $\text{W}{xi}$, $\text{W}{hi}$, $\text{W}{xf}$, $\text{W}{hf}$, $\text{W}{xo}$, and $\text{W}{ho}$, depicted in Figure 8, manage information from inputs ($x_{t}$) and history ($h_{t-1}$$C_{t-1}$) in a convolutional manner.
Metrics U-NO F-FNO MWT FNO
RMSE 0.3675 0.2280 0.3494 0.1878
MAE 0.1465 0.1350 0.2228 0.1634
Table 4: The Result of U-NO, F-FNO, MWT and FNO in the FN Case.
C: Baseline Models
In order to compare and evaluate the performance of our proposed method, we have trained multiple state-of-the-art (SOTA) baseline models as well as classical models, and compared them with our model. The introductions to each baseline model are presented below, while the training details are outlined in subsequent sections D: Training Details.
Fourier Neural Operator (FNO) (li2020fourier). FNO is a method that combines Fourier transforms with neural networks. This approach comprises two main components. The first component involves performing Fourier transforms on the system state quantities, learning certain information in the frequency domain, and then applying an inverse transform. The second component utilizes convolutions to process the system state quantities, complementing the information not captured during the frequency domain learning. The combination of these two components serves as the final result. We make two FNO models sharing the same architecture, which training in micro and macro scale datasets respectively and inferring in micro and macro time intervals. In the same model, we conducted training separately for two types of data, resulting in FNO and FNO-coarse.
ConvLSTM (vlachas2022multiscale). ConvLSTM is a specialized neural network architecture that combines convolutional and LSTM layers to effectively model spatial and temporal dependencies in sequential data. We use AE-ConvLSTM as our base model.
PeRCNN (Rao_2023). PeRCNN represents a physics-informed learning methodology, embedding physical laws directly into the neural network architecture. It incorporates multiple parallel convolutional neural networks (CNNs), leveraging the simulation of polynomial equations through feature map multiplication. By doing so, PeRCNN augments the model’s extrapolation and generalization capabilities.
D: Training Details
@@
E: Evaluation Metrics
We have adopted some classical evaluation metrics such as RMSE, MAE and HCT. Root Mean Square Error (RMSE) quantifies the average error magnitude between estimated and actual values, serving as a gauge of the model’s precision. Conversely, Mean Absolute Error (MAE) assesses the average absolute disparity between anticipated and observed values, denoting the true scale of discrepancies.
@@
RMSE (Root Mean Square Error): $\quad\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (8)
@@
[…]
In the above equations8, $n$ represents the number of trajectories, $y_{i}$ represents the true value, and $\hat{y}_{i}$ represents the predicted value of the model. The PCC is the Pearson correlation coefficient, which is a statistical metric used to measure the linear correlation between two variables.
F: Dataset Informations
@@
Model Running Time Parameter Size GPU Memory
PIMRL 9 s 3.33 M 1728 M
U-NO 7 s 15.29 M 2320 M
MWT 12 s 0.09 M 1732 M
FNO 2 s 8.39 M 1580 M
ConvLSTM 5 s 3.32 M 1708 M
Table 5: Running Time, Parameter Size, and GPU Memory for PIMRL, U-NO, MWT, FNO, and ConvLSTM
@@
Table 6: Summary of experimental settings for different cases.(The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training)
@@
$\frac{\partial u}{\partial t}=-u\frac{\partial u}{\partial x}-\frac{\partial^{3}u}{\partial x^{3}}$ (9)
@@
$\frac{\partial u}{\partial t}$ $=-uu_{x}-vu_{y}+\nu\nabla^{2}u,$ (10)
$\frac{\partial v}{\partial t}$ $=-uv_{x}-vv_{y}+\nu\nabla^{2}v.$ (11)
@@
$\frac{\partial u}{\partial t}$ $=\mu_{u}\nabla^{2}u+u-u^{3}-v+\alpha,$ (12)
$\frac{\partial v}{\partial t}$ $=\mu_{v}\nabla^{2}v+(u-v)\beta.$ (13)
@@
$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+F(1-u),$ (14)
$\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(F+k)v.$ (15)
@@
In the three-dimensional case, we have the parameters: $D_{u}=0.2$, $D_{v}=0.1$, $F=0.025$, and $k=0.055$. We got 3 trajectories for training, 1 trajectory for validation and 2 trajectories for testing, where $\Delta t=15\delta t$ and $\delta t=0.25s$.
G: Supplement results